OpenStudy (anonymous):

every evening jenna empties her pockets and puts her change in a jar. At the end of the week she counts her money. One week she had 38 coins, all of them dimes and quarters. When she added them up she had a total of $6.95 let d=number of dimes q= the number of quarters how many dimes did jenna have?

OpenStudy (tylerd):

6.95=0.10d+0.25q d+q=38

OpenStudy (tylerd):

q=38-d 6.95=0.10d+0.25(38-d)

OpenStudy (tylerd):

6.95=0.10d-0.25d+9.5

OpenStudy (tylerd):

solve for d then 38-d=q